OpenStudy (sarahc):

How many milliliters of sodium metal, with a density of 0.97 g/mL, would be needed to produce 43.6 grams of sodium hydroxide in the single replacement reaction below? Be sure to show the work that you did to solve this problem. unbalanced equation: Na + H2O NaOH + H2

OpenStudy (anonymous):

first, we need to balance the equation because it relates the number of sodium atoms consumed to the number of NaOH ion pairs produced. would you like to try?

OpenStudy (sarahc):

46×43.6÷2=1002.8 g

OpenStudy (sarahc):

=1002.8/0.97=1033.8 ml ??

OpenStudy (sarahc):

oh, 2Na+2H2O---2NaOH+H2

OpenStudy (anonymous):

i agree with your balanced equation, but not the mass of sodium

OpenStudy (sarahc):

Its the grams of sodium...I thought thats how you do it...:/ Do u mind showing me how?

OpenStudy (anonymous):

i would find the number of NaOH pairs like this: divide 43.6 grams by the formula weight (how much does 1 mol of NaOH pairs weigh?) then i would convert it to the number of Na atoms: number of Na = (number of NaOH pairs) * (2 Na / 2 NaOH pairs) the 2's are the numbers in front of the things in the balanced equation.